# Kovács’ conjecture on characterization of projective space and hyperquadrics

### v1

@@ Abstract.
We prove Kovács’ conjecture that claims that if the $p^{th}$ exterior power of the tangent bundle of a smooth complex projective variety contains the $p^{th}$ exterior power of an ample vector bundle then the variety is either projective space or the $p$-dimensional quadric hypersurface. This provides a common generalization of Mori, Wahl, Cho–Sato, Andreatta–Wiśniewski, Kobayashi–Ochiai, and Araujo–Druel–Kovács type characterizations of such varieties.
@@ 1. Introduction
Conjecture I was proven in [Ross11] for varieties with Picard number $1$ ([Ross11][Theorem 1.1]) and for $n=2$ ([Ross11][Corollary 1.10]). Furthermore, [Liu23][Theorem 1.15] proves Conjecture I for $n\geq 3$, $p=2$ and $\binom{r}{2}\geq n-1$.
@@ 1.1. Existing results and background
Using the theory of variety of minimal rational tangents, Araujo provided a unified geometric approach to Theorems 1.2– 1.4 in [Ara06]. The following characterization for projective spaces and higher dimensional quadric hypersurfaces was proven in 1973 by Kobayashi and Ochiai.
@@
Despite their similar nature, no one of Theorems 1.2– 1.7 imply all the other. From this perspective, Conjecture I is interesting, since it implies all of the above: Theorem 1.2 follows by taking $p=1$, $\mathcal{E}=T_{X}$, Theorem 1.3 by p=1,r=1p=1,\ r=1, Theorem 1.4 by $p=1$, Theorem 1.6 (for characteristic 0) by $p=2$ and $\mathcal{E}=T_{X}$, Theorem 1.7 by taking $\mathcal{E}=\mathcal{L}^{\oplus p}$, and Theorem 1.5 by taking $\mathcal{E}=\mathcal{L}^{\oplus n}$ and $\mathcal{L}^{\oplus n-1}\oplus\mathcal{L}^{\otimes 2}$. Indeed, in [Ross11], Ross proves Conjecture I for varieties with Picard number $1$ and also for $n=2$. More recently in [Liu23], Liu proved the conjecture in higher dimensions for $p=2$ under the rank constraint $\binom{r}{2}\geq n-1$.
@@
Recent results have focused on weakening the positivity assumption from ample to nef or strictly nef, at the cost of rigidity of the structure. Notably, Li–Ou–Yang [LOY19] obtained $X\cong\mathbb{P}^{n}$ or $Q_{n}$, assuming $T_{X}$ is strictly nef, or $n\geq 3$ and $\bigwedge^{2}T_{X}$ is strictly nef.
@@ 2.1. Characterizing the variety
In this section, we prove a weaker version of Theorem 1.1, namely, we obtain a weaker characterization of $\mathcal{E}$, when $X\cong\mathbb{P}^{n}$. Let $X$ be a smooth complex projective variety of dimension $n$ and let $\mathcal{E}$ be ample vector bundle on $X$ of rank $r$. For any positive integer integer $1\leq p\leq r$. Set
$f:=\operatorname{rank}\left(\bigwedge^{p}\mathcal{E}\right)=\binom{r}{p},\qquad a:=\binom{r-1}{p-1}$.
We have the binomial identity
$p\binom{r}{p}=r\binom{r-1}{p-1},\text{ i.e.},\ pf=ra.$ (2.1)
@@
For any vector bundle $\mathcal{E}$ of rank $r$ and $p\leq r$, $\det\left(\bigwedge^{p}\mathcal{E}\right)\ \cong\ \det(\mathcal{E})^{\otimes a}$, where $a=\binom{r-1}{p-1}$.
@@
Choose a trivializing open cover $X=\bigcup_{i\in I}\mathcal{U}_{i}$ for $\mathcal{E}$, i.e., $\mathcal{E}|_{\mathcal{U}_{i}}\cong\mathcal{O}_{\mathcal{U}_{i}}^{\oplus r}$ and given by transition fucntions $\{\varepsilon_{ij}\in\operatorname{GL}_{r}(\mathcal{O}_{X}(\mathcal{U}_{ij}))\mid i\neq j\in I\}$. Then the bundle $\bigwedge^{p}\mathcal{E}$ is given by the same trivializing cover and transition functions $\{\wedge^{p}\varepsilon_{ij}\in\operatorname{GL}_{\binom{r}{p}}(\mathcal{O}_{X}(\mathcal{U}_{ij}))\mid i\neq j\in I\}$, and the line bundle $\det\bigwedge^{p}\mathcal{E}$ is given by the transition functions $\{\det(\wedge^{p}\varepsilon_{ij})\in\mathcal{O}_{X}(\mathcal{U}_{ij})^{\times}\mid i\neq j\in I\}$. By the classical Sylvester-Franke theorem in multilinear algebra [Price47][§6], we have at the level of transition functions, $\det(\wedge^{p}\varepsilon_{ij})=(\det\varepsilon_{ij})^{a}$, where $a=\binom{r-1}{p-1}$. The claim then follows. ∎
@@
If there is a subbundle inclusion $\bigwedge^{p}\mathcal{E}\subseteq\bigwedge^{p}T_{X}$ for some positive integer $p\leq r$, then
$H^{0}\left(X,\ T_{X}^{\otimes pf}\otimes\det(\mathcal{E})^{-a}\right)\neq 0.$
@@
Thus, we obtain a nonzero section in $H^{0}\bigl(X,\ T_{X}^{\otimes pf}\otimes\det(\bigwedge^{p}\mathcal{E})^{-1}\bigr)$. By Lemma 2.1, the proof is complete. ∎
@@
Now by Druel-Paris Theorem 1.8. Then we have two possibilities:
@@ 2.2. Characterizing $\mathcal{E}$ when $X\cong\mathbb{P}^{n}$
In this section we characterize ample vector bundles $\mathcal{E}$ of rank $r$ on $\mathbb{P}^{n}$ such that
  1. (i)
    $\bigwedge^{p}\mathcal{E}\subseteq\bigwedge^{p}T_{\mathbb{P}^{n}}$ and,
  2. (ii)
    $\det\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $r\leq l\leq\frac{r(n+1)}{n}$.
Clearly, we have the constraints $0<p\leq n$ and $0<r\leq n$. First note the following well-known splitting type of the tangent bundle $T_{\mathbb{P}^{n}}$ on any line $L\subset\mathbb{P}^{n}$ (see for example, [OSS88][Example, Pg 14]).
@@
We now find the splitting type of ample vector bundle $\mathcal{E}$ on $\mathbb{P}^{n}$ satisfying condition $(i)$ and $(ii)$.
@@
Let $L\cong\mathbb{P}^{1}$ be any line. Let $\mathcal{E}$ be an ample vector bundle of rank $r$ on $\mathbb{P}^{n}$ satisfying (i)(i)$(ii)$ above. Then $l:=\deg\det\mathcal{E}=r$ or $r+1$ and:
@@
Proposition 2.7 gives an independent proof of $l=r$ or $r+1$, which we also obtained in Theorem 2.3 from Druel-Paris’s bounds in Theorem 1.8.
@@
Let $\mathcal{E}$ be an ample vector bundle of rank $r$ on $\mathbb{P}^{n}$ such that $(a)$ $\bigwedge^{p}\mathcal{E}\subseteq\bigwedge^{p}T_{\mathbb{P}^{n}}$ for some positive integer $p\leq r$ and, $(b)$ $\det\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $r\leq l\leq r(n+1)/n$. Then $r\leq n$ and:
@@
Case 1: $r<n$. If $r<n$, then $r+1>r(n+1)/n$, and so $r\leq l\leq r(n+1)/n$ implies $l=r$. By Proposition 2.7, the bundle $\mathcal{E}$ is uniform with splitting type $(1,\dots,1)$. Applying Theorem 2.9 (with $r<n$) shows that $\mathcal{E}$ splits as a sum of line bundles, say $\mathcal{E}\cong\bigoplus_{i=1}^{r}\mathcal{O}_{\mathbb{P}^{n}}(a_{i})$. Restricting to any line $L$ and using Proposition 2.7, each $a_{i}$ must equal $1$, hence $\mathcal{E}\cong\mathcal{O}(1)^{\oplus r}$.
Case 2: $r=n$ and $l=n$. Then Proposition 2.7 gives $\mathcal{E}|_{L}\cong\mathcal{O}_{L}(1)^{\oplus n}$ on every line $L$. By Theorem 2.9, $\mathcal{E}$ is either a sum of line bundles or a twist of $T_{\mathbb{P}^{n}}$ or $\Omega^{1}_{\mathbb{P}^{n}}$. But the latter two possibilities are excluded by Lemma 2.5: $T_{\mathbb{P}^{n}}|_{L}\ \cong\ \mathcal{O}_{L}(2)\ \oplus\ \mathcal{O}_{L}(1)^{\oplus(n-1)}$ and $\Omega^{1}_{\mathbb{P}^{n}}|_{L}\ \cong\ \mathcal{O}_{L}(-2)\ \oplus\ \mathcal{O}_{L}(-1)^{\oplus(n-1)}$. Thus $\mathcal{E}\cong\oplus_{i}\mathcal{O}(a_{i})$, and as in Case 1, we have $\mathcal{E}\cong\mathcal{O}(1)^{\oplus n}$.
Case 3: $r=n$ and $l=n+1$. By Proposition 2.7, $\mathcal{E}$ is uniform with splitting type $(2,1,\dots,1)$ on every line. By Theorem 2.9 it is either a direct sum of line bundles or a twist of $T_{\mathbb{P}^{n}}$ or $\Omega^{1}_{\mathbb{P}^{n}}$.

### v2

@@ Abstract.
We prove Kovács’ conjecture that claims that if the $p^{th}$ exterior power of the tangent bundle of a smooth complex projective variety contains the $p^{th}$ exterior power of an ample vector bundle then the variety is either projective space or the $p$-dimensional quadric hypersurface. We also prove a similar characterization involving symmetric powers instead of exterior powers. This provides a common generalization of Mori, Wahl, Cho–Sato, Andreatta–Wiśniewski, Kobayashi–Ochiai, and Araujo–Druel–Kovács type characterizations of such varieties.
@@ 1. Introduction
Conjecture I was proven in [Ross11] for varieties with Picard number $1$ ([Ross11][Theorem 1.1]) and for $n=2$ ([Ross11][Corollary 1.10]). Furthermore, [Liu23][Theorem 1.15] proves Conjecture I for $n\geq 3$, $p=2$ and $\binom{r}{2}\geq n-1$. Using similar technique as for proof of Theorem 1.1, we also prove the following characterization using symmetric powers of vector bundles.
Theorem 1.2.
Let $X$ be a smooth complex projective variety of dimension $n$. If there exists an ample vector bundle $\mathcal{E}$ of rank $r$ on $X$ and a positive integer $p$ such that $\operatorname{Sym}^{p}\mathcal{E}\subseteq\operatorname{Sym}^{p}T_{X}$, then either $X\cong\mathbb{P}^{n}$, or $(X,\mathcal{E})\cong(Q_{n},\mathcal{O}_{Q_{n}}(1)^{\oplus r})$ where $Q_{n}\subseteq\mathbb{P}^{n+1}$ is a smooth quadric hypersurface in $\mathbb{P}^{n+1}$. Furthermore, if $X\cong\mathbb{P}^{n}$, then $\det\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $l=r$ or $r+1$ and:
  1. (1)
    If $r<n$, then $\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(1)^{\oplus r}$.
  2. (2)
  3. (3)
    If $r=n$ and $l=n+1$, then $\mathcal{E}\cong T_{\mathbb{P}^{n}}$.
@@ 1.1. Existing results and background
Using the theory of variety of minimal rational tangents, Araujo provided a unified geometric approach to Theorems 1.3– 1.5 in [Ara06]. The following characterization for projective spaces and higher dimensional quadric hypersurfaces was proven in 1973 by Kobayashi and Ochiai.
@@
Despite their similar nature, no one of Theorems 1.3– 1.8 imply all the other. From this perspective, Conjecture I is interesting, since it implies all of the above: Theorem 1.3 follows by taking $p=1$, $\mathcal{E}=T_{X}$, Theorem 1.4 by p=1,r=1p=1,\ r=1, Theorem 1.5 by $p=1$, Theorem 1.7 (for characteristic 0) by $p=2$ and $\mathcal{E}=T_{X}$, Theorem 1.8 by taking $\mathcal{E}=\mathcal{L}^{\oplus p}$, and Theorem 1.6 by taking $\mathcal{E}=\mathcal{L}^{\oplus n}$ and $\mathcal{L}^{\oplus n-1}\oplus\mathcal{L}^{\otimes 2}$. Indeed, in [Ross11], Ross proves Conjecture I for varieties with Picard number $1$ and also for $n=2$. More recently in [Liu23], Liu proved the conjecture in higher dimensions for $p=2$ under the rank constraint $\binom{r}{2}\geq n-1$.
@@
Recent results have focused on weakening the positivity assumption from ample to nef or strictly nef, at the cost of rigidity of the structure. Notably, Li–Ou–Yang [LOY19] obtained $X\cong\mathbb{P}^{n}$ or $Q_{n}$, assuming $T_{X}$ is strictly nef, or $n\geq 3$ and $\bigwedge^{2}T_{X}$ is strictly nef. Recently, Shao [Shao24] has also conjectured a characterization of projective space and hyperquadrics using cohomology of twists of symmetric powers of tangent bundle and has proven it for varieties with Picard number $1$.
@@ 2.1. Characterizing the variety
In this section, we prove weaker versions of Theorem 1.1 and Theorem 1.2, namely, we obtain a weaker characterization of $\mathcal{E}$, when $X\cong\mathbb{P}^{n}$. Let $X$ be a smooth complex projective variety of dimension $n$ and let $\mathcal{E}$ be ample vector bundle on $X$ of rank $r$. For any positive integer integer $1\leq p\leq r$. For the remainder of the paper, set
$f:=\operatorname{rank}\left(\bigwedge^{p}\mathcal{E}\right)=\binom{r}{p},\ a:=\binom{r-1}{p-1};\qquad f^{\prime}:=\operatorname{rank}\left(\operatorname{Sym}^{p}\mathcal{E}\right)=\binom{p+r-1}{p},\ a^{\prime}:=\binom{p+r-1}{r}$.
We have the binomial identities
$p\binom{r}{p}=r\binom{r-1}{p-1},\text{ i.e.},\ pf=ra;\qquad p\binom{p+r-1}{p}=r\binom{p+r-1}{r},\text{ i.e.},\ pf^{\prime}=ra^{\prime}.$ (2.1)
@@
For any vector bundle $\mathcal{E}$ of rank $r$ and $p\leq r$, we have $\det\left(\bigwedge^{p}\mathcal{E}\right)\ \cong\ \det(\mathcal{E})^{\otimes a}$, where $a=\binom{r-1}{p-1}$ and $\det\left(\operatorname{Sym}^{p}\mathcal{E}\right)\ \cong\ \det(\mathcal{E})^{\otimes a^{\prime}}$, where $a^{\prime}=\binom{p+r-1}{r}$.
@@
Note that $\det(\bigwedge^{p}(-))$ and $\det(\operatorname{Sym}^{p}(-))$ are $1$-dimensional representations of the general linear group and hence $\det(\bigwedge^{p}(-))=\det(-)^{\otimes b}$ and $\det(\operatorname{Sym}^{p}(-))=\det(-)^{\otimes b^{\prime}}$ for some integers $b,b^{\prime}$. Fiberwise, a scalar $\lambda$ acts on $\bigwedge^{p}\mathcal{E}$ and $\operatorname{Sym}^{p}\mathcal{E}$ as multiplication by $\lambda^{p}$. Therefore, fiberwise $\lambda$ acts on $\det\bigwedge^{p}\mathcal{E}$ by multiplication by $\lambda^{pf}=\lambda^{ra}$ by (2.1), which implies $b=a$. Similarly, fiberwise $\lambda$ acts on $\det\operatorname{Sym}^{p}\mathcal{E}$ by multiplication by $\lambda^{pf^{\prime}}=\lambda^{ra^{\prime}}$ by (2.1), which implies $b^{\prime}=a^{\prime}$.
We note that alternatively, one can take a trivializing cover for $\mathcal{E}$ on $X$ and apply the classical Sylvester-Franke theorem in multilinear algebra [Price47][§6] to $\det(\bigwedge^{p}\mathcal{E})$ and $\det(\operatorname{Sym}^{p}\mathcal{E})$ at the level of transition functions to obtain the desired identity. ∎
@@
If there is a subbundle inclusion $\bigwedge^{p}\mathcal{E}\subseteq\bigwedge^{p}T_{X}$ for any positive integer $p\leq r$, then
$H^{0}\left(X,\ T_{X}^{\otimes pf}\otimes\det(\mathcal{E})^{-a}\right)\neq 0.$ (2.2)
Similarly, if there is a subbundle inclusion $\operatorname{Sym}^{p}\mathcal{E}\subseteq\operatorname{Sym}^{p}T_{X}$ for any positive integer $p$, then
$H^{0}\left(X,\ T_{X}^{\otimes pf^{\prime}}\otimes\det(\mathcal{E})^{-a^{\prime}}\right)\neq 0.$ (2.3)
@@
Thus, we obtain a nonzero section in $H^{0}\bigl(X,\ T_{X}^{\otimes pf}\otimes\det(\bigwedge^{p}\mathcal{E})^{-1}\bigr)$ and proves (2.2) by Lemma 2.1.
Similarly, if $\iota^{\prime}:\operatorname{Sym}^{p}\mathcal{E}\hookrightarrow\operatorname{Sym}^{p}T_{X}$ is the inclusion, taking the induced map on $(f^{\prime})^{th}$ exterior powers gives the injection $\wedge^{f^{\prime}}\iota^{\prime}:\det(\operatorname{Sym}^{p}\mathcal{E})\hookrightarrow\bigwedge^{f^{\prime}}(\operatorname{Sym}^{p}T_{X})$. Using the antisymmetrization map, we have an injective map $\alpha^{\prime}_{f^{\prime}}:\bigwedge^{f^{\prime}}(\operatorname{Sym}^{p}T_{X})\hookrightarrow(\operatorname{Sym}^{p}T_{X})^{\otimes f^{\prime}}$. Analogous to the antisymmetrization map, for any vector bundle $\mathcal{V}$ of rank $r$, there are canonical symmetrization maps $\gamma_{k}:\operatorname{Sym}^{k}\mathcal{V}\rightarrow\mathcal{V}^{\otimes k}$ for any $k\geq 1$, given on sections by $\gamma_{k}(v_{1}\cdot v_{2}\cdots v_{k})=\sum_{\sigma\in\Sigma_{k}}v_{\sigma(1)}\otimes v_{\sigma(2)}\otimes\cdots v_{\sigma(k)}$. When $k!$ is invertible in the base field, the symmetrization map is injective. The $p^{th}$ symmetrization map for $\mathcal{V}=T_{X}$ therefore gives an injective map $\gamma_{p}:\operatorname{Sym}^{p}T_{X}\rightarrow T_{X}^{\otimes p}$, which therefore induces an injective map $\gamma_{p}^{\otimes f^{\prime}}:(\operatorname{Sym}^{p}T_{x})^{\otimes f^{\prime}}\hookrightarrow T_{X}^{\otimes pf^{\prime}}$. Thus, the composition
$\det\left(\operatorname{Sym}^{p}\mathcal{E}\right)\xrightarrow{\wedge^{f^{\prime}}\iota^{\prime}}\bigwedge^{f^{\prime}}\left(\operatorname{Sym}^{p}T_{X}\right)\xrightarrow{\alpha^{\prime}_{f^{\prime}}}\left(\operatorname{Sym}^{p}T_{X}\right)^{\otimes f^{\prime}}\xrightarrow{\gamma_{p}^{\otimes f^{\prime}}}T_{X}^{\otimes pf^{\prime}}$ (2.8)
yields a non-zero section in $H^{0}\bigl(X,\ T_{X}^{\otimes pf^{\prime}}\otimes\det(\operatorname{Sym}^{p}\mathcal{E})^{-1}\bigr)$ and proves (2.3) by Lemma 2.1. ∎
@@
Now by Druel-Paris Theorem 1.9, we have two possibilities:
@@
Theorem 2.5.
Assume $\operatorname{Sym}^{p}\mathcal{E}\subseteq\operatorname{Sym}^{p}T_{X}$ for an ample rank $r$ vector bundle $\mathcal{E}$ on $X$, and some positive integer $p$. Then $(X,\det\mathcal{E})\cong(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}^{n}}(l))$ with $r\leq l\leq r(n+1)/n$, or $(X,\mathcal{E})\cong(Q_{n},\mathcal{O}_{Q_{n}}(1)^{\oplus r})$ where $Q_{n}\subseteq\mathbb{P}^{n+1}$ is a smooth hyperquadric.
Proof.
Consider the ample vector bundle $\mathcal{F}:=\mathcal{E}^{\oplus a^{\prime}}$ on $X$. Then $\operatorname{rank}(\mathcal{F})=ra^{\prime}$ and $\det(\mathcal{F})=\det(\mathcal{E})^{\otimes a^{\prime}}$. But we have $pf^{\prime}=ra^{\prime}=\operatorname{rank}(\mathcal{F})$ by (2.1), and Lemma 2.2 gives us
$H^{0}\left(X,\ T_{X}^{\otimes pf^{\prime}}\otimes\det(\mathcal{F})^{-1}\right)\neq 0.$
Thus, Druel-Paris Theorem 1.9, gives us the two possibilities:
  1. (1)
    $(X,\det\mathcal{F})\cong(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}^{n}}(l))$ with $\operatorname{rank}\mathcal{F}\leq l\leq\operatorname{rank}\mathcal{F}(n+1)/n$. But $\det\mathcal{F}=\det\mathcal{E}^{\oplus a^{\prime}}=(\det\mathcal{E})^{\otimes a^{\prime}}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$, with $ra^{\prime}\leq l\leq ra^{\prime}(n+1)/n$. Therefore, $\det\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $r\leq l\leq r(n+1)/n$ where $r=\operatorname{rank}\mathcal{E}$.
  2. (2)
    $(X,\mathcal{F})\cong(Q_{n},\mathcal{O}_{Q_{n}}(1)^{\oplus k})$ where $k=2i+nj$ with $i,j\gg 0$. But $k=\operatorname{rank}\mathcal{F}=ra^{\prime}$, where $r=\operatorname{rank}\mathcal{E}$. Thus, $\mathcal{E}=\mathcal{O}_{Q_{n}}(1)^{\oplus r}$.
@@ 2.2. Characterizing $\mathcal{E}$ when $X\cong\mathbb{P}^{n}$
In this section we characterize ample vector bundles $\mathcal{E}$ of rank $r$ on $\mathbb{P}^{n}$ satisfying either of the following two conditions:
  • Condition A: $\bigwedge^{p}\mathcal{E}\subseteq\bigwedge^{p}T_{\mathbb{P}^{n}}$ and $\det\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $r\leq l\leq\frac{r(n+1)}{n}$ or,
  • Condition B: $\operatorname{Sym}^{p}\mathcal{E}\subseteq\operatorname{Sym}^{p}T_{\mathbb{P}^{n}}$ and $\det\mathcal{E}\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $r\leq l\leq\frac{r(n+1)}{n}$.
Clearly, we have $0<r\leq n$ in either situations as $\operatorname{rank}(\bigwedge^{p}\mathcal{V})$ and $\operatorname{rank}(\operatorname{Sym}^{p}\mathcal{V})$ are both increasing functions of $\operatorname{rank}(\mathcal{V})$. In Condition A, we also have the constraint $0<p\leq n$, whereas in Condition B, $p$ is any positive integer. First note the following well-known splitting type of the tangent bundle $T_{\mathbb{P}^{n}}$ on any line $L\subset\mathbb{P}^{n}$ (see for example, [OSS88][Example, Pg 14]).
@@
Analogously, we have the following for symmetric powers of $T_{\mathbb{P}^{n}}$.
Lemma 2.8.
For $p>0$ and a line $L\cong\mathbb{P}^{1}$ in $\mathbb{P}^{n}$, one has
$\left(\operatorname{Sym}^{p}T_{\mathbb{P}^{n}}\right)|_{L}\cong\operatorname{Sym}^{p}(T_{\mathbb{P}^{n}}|_{L})\ \cong\ \bigoplus_{\begin{subarray}{c}m_{1}+m_{2}+\dots+m_{n}=p\\ m_{i}\geq 0\end{subarray}}\mathcal{O}_{L}(2m_{1}+m_{2}+\dots+m_{n}).$
In particular, every line bundle in the splitting type of $\operatorname{Sym}^{p}T_{\mathbb{P}^{n}}$ has degree $p+m_{1}$ for some $0\leq m_{1}\leq p$, and the maximal degree occurring is $2p$, with multiplicity $1$.
Proof.
By Lemma 2.6, $T_{\mathbb{P}^{n}}|_{L}\cong\mathcal{O}_{L}(2)\oplus\mathcal{O}_{L}(1)^{\oplus(n-1)}$. The splitting type of $\operatorname{Sym}^{p}T_{\mathbb{P}^{n}}$ follows from noting that direct sum and symmetric powers commute. Clearly the maximal degree line bundle in the splitting can only occur when $m_{1}=p$ and $m_{i}=0$ for all $i\geq 2$, which gives degree $2p$ with multiplicity $1$. ∎
We now find the splitting type of ample vector bundle $\mathcal{E}$ on $\mathbb{P}^{n}$ satisfying Condition A.
@@
Let $L\cong\mathbb{P}^{1}$ be any line. Let $\mathcal{E}$ be an ample vector bundle of rank $r$ on $\mathbb{P}^{n}$ satisfying Condition A above. Then $l:=\deg\det\mathcal{E}=r$ or $r+1$ and:
@@
A very similar analysis yields the splitting type of an ample vector bundle $\mathcal{E}$ on $\mathbb{P}^{n}$ satisfying Condition B.
Proposition 2.10.
Let $L\cong\mathbb{P}^{1}$ be any line. Let $\mathcal{E}$ be an ample vector bundle of rank $r$ on $\mathbb{P}^{n}$ satisfying Condition B above. Then $l:=\deg\det\mathcal{E}=r$ or $r+1$ and:
  1. (1)
  2. (2)
Proof.
As in the above proof, $\mathcal{E}|_{L}\cong\bigoplus_{i=1}^{r}\mathcal{O}_{L}(a_{i})$ with $a_{1}\geq a_{2}\geq\cdots\geq a_{r}\geq 1$. Restrict the given injection $\iota^{\prime}:\operatorname{Sym}^{p}\mathcal{E}\hookrightarrow\operatorname{Sym}^{p}T_{\mathbb{P}^{n}}$ to a line $L\subset\mathbb{P}^{n}$ to obtain
$\iota^{\prime}|_{L}:\operatorname{Sym}^{p}(\mathcal{E}|_{L})\ \hookrightarrow\ \operatorname{Sym}^{p}(T_{\mathbb{P}^{n}}|_{L}).$
By Lemma 2.8, the target is a direct sum of copies of $\mathcal{O}_{L}(m_{1}+p)$, where $0\leq m_{i}\leq p$. On the other hand, $\operatorname{Sym}^{p}(\mathcal{E}|_{L})$ splits as a direct sum of line bundles $\mathcal{O}_{L}(a_{1}m_{1}+a_{2}m_{2}+\dots+a_{r}m_{r})$, where $m_{i}\geq 0$ and $m_{1}+\dots+m_{r}=p$. If $a_{1}>2$, then one of the summands in $\operatorname{Sym}^{p}(\mathcal{E}|_{L})$ would have degree $pa_{1}>2p$, which is the maximal degree among line bundles occurring in $\operatorname{Sym}^{p}(T_{\mathbb{P}^{n}}|_{L})$. This would violate the inclusion $\iota^{\prime}|_{L}:\operatorname{Sym}^{p}(\mathcal{E}|_{L})\ \hookrightarrow\ \operatorname{Sym}^{p}(T_{\mathbb{P}^{n}}|_{L})$ and thus, $a_{1}\leq 2$. Similarly if $a_{1}=a_{2}=2$, then the multiplicity of $\mathcal{O}_{L}(2p)$ in $\operatorname{Sym}^{p}(\mathcal{E}|_{L})$ would be the number of solutions $m_{1}+m_{2}=p$, with $m_{i}\geq 0$, which is greater than $1$. Since $\mathcal{O}_{L}(2p)$ occurs with multiplicity $1$ in $\operatorname{Sym}^{p}(T_{\mathbb{P}^{n}}|_{L})$, we see that $a_{2}=2$ violates the inclusion $\iota^{\prime}|_{L}:\operatorname{Sym}^{p}(\mathcal{E}|_{L})\ \hookrightarrow\ \operatorname{Sym}^{p}(T_{\mathbb{P}^{n}}|_{L})$. Thus $a_{1}\in\{1,2\}$ and all other $a_{i}=1$ for $2\leq i\leq r$, proving the stated two possible splitting types on a line. Finally, as in the proof of Proposition 2.9, $\det(\mathcal{E})\cong\mathcal{O}_{\mathbb{P}^{n}}(\sum_{i=1}^{r}a_{i})\cong\mathcal{O}_{\mathbb{P}^{n}}(l)$ with $l$ equal to $r$ or $r+1$, depending on whether $(a_{1},\dots,a_{r})=(1,1,\dots,1)$ or $(2,1,1\dots,1)$ respectively. This also shows that $\mathcal{E}$ is uniform. ∎
@@
Propositions 2.9 and  2.10 give independent proofs of $l=r$ or $r+1$, which we also obtained in Theorems 2.3 and  2.5 from Druel-Paris’s bounds in Theorem 1.9.
@@
Let $\mathcal{E}$ be an ample vector bundle of rank $r$ on $\mathbb{P}^{n}$ satisfying either Condition (A) or (B). Then $r\leq n$ and:
@@
  4. (4)
    If $r=n$, $l=n+1$ and $\mathcal{E}$ satisfies Condition B, then $\mathcal{E}\cong T_{\mathbb{P}^{n}}$.
@@
Case 1: $r<n$. If $r<n$, then $r+1>r(n+1)/n$, and so $r\leq l\leq r(n+1)/n$ implies $l=r$. By Proposition 2.9 and 2.10, the bundle $\mathcal{E}$ is uniform with splitting type $(1,\dots,1)$ irrespective of whether $\mathcal{E}$ satisfies Condition A or B. Applying Theorem 2.12 (with $r<n$) shows that $\mathcal{E}$ splits as a sum of line bundles, say $\mathcal{E}\cong\bigoplus_{i=1}^{r}\mathcal{O}_{\mathbb{P}^{n}}(a_{i})$. Restricting to any line $L$ and using Proposition 2.9, each $a_{i}$ must equal $1$, hence $\mathcal{E}\cong\mathcal{O}(1)^{\oplus r}$.
Case 2: $r=n$ and $l=n$. Then Proposition 2.9 and 2.10 give $\mathcal{E}|_{L}\cong\mathcal{O}_{L}(1)^{\oplus n}$ on every line $L$, irrespective of whether $\mathcal{E}$ satisfies Condition A or B. By Theorem 2.12, $\mathcal{E}$ is either a sum of line bundles or a twist of $T_{\mathbb{P}^{n}}$ or $\Omega^{1}_{\mathbb{P}^{n}}$. But the latter two possibilities are excluded by Lemma 2.6: $T_{\mathbb{P}^{n}}|_{L}\ \cong\ \mathcal{O}_{L}(2)\ \oplus\ \mathcal{O}_{L}(1)^{\oplus(n-1)}$ and $\Omega^{1}_{\mathbb{P}^{n}}|_{L}\ \cong\ \mathcal{O}_{L}(-2)\ \oplus\ \mathcal{O}_{L}(-1)^{\oplus(n-1)}$. Thus $\mathcal{E}\cong\oplus_{i}\mathcal{O}(a_{i})$, and as in Case 1, we have $\mathcal{E}\cong\mathcal{O}(1)^{\oplus n}$.
Case 3: $r=n$, $l=n+1$ and $\mathcal{E}$ satisfies Condition A. By Proposition 2.9, $\mathcal{E}$ is uniform with splitting type $(2,1,\dots,1)$ on every line. By Theorem 2.12 it is either a direct sum of line bundles or a twist of $T_{\mathbb{P}^{n}}$ or $\Omega^{1}_{\mathbb{P}^{n}}$.